OpenStudy (anonymous):

I don't know if this has been posted, but recently it was announced that a couple of physicists were able to design and experiment in which matter (in the form of an electron and a positron) would be created from two photons of light.

OpenStudy (anonymous):

This experiment would prove a theory developed in the 1930's by Breit and Wheeler. So the question is: how the heck do you get matter from energy?

OpenStudy (anonymous):

Einstein gave part of that answer with his matter-energy equivalency statement: \[E=mc ^{2}\]Max Planck gave the rest of the answer with his statement of the quantized energy of a photon:\[E=h \nu\]

OpenStudy (anonymous):

If we set those equations equal to each other we get:\[mc ^{2}=h \nu\]m is mass, and in the case of this experiment is equal to the sum of the mass of an electron and a positron...both of which weigh the same: 9.11*10^-31kg.

OpenStudy (anonymous):

The conversion requires two photons, so the equation now looks like this:\[2m _{e}c ^{2}=h(2 \nu )\]which just simplifies to:\[m _{e}c ^{2}=h \nu\]Solving for ν we get:\[\nu =\frac{m _{e}c ^{2} }{ h }=\frac{ \left( 9.11*10^{-31}kg \right)\left(2.99*10^{8}m/s\right)^{2} }{6.626*10^{-34}J∙s }\]

OpenStudy (anonymous):

This yields a photon frequency of 1.2*10^20 Hz, a photon in the gamma ray range of frequencies. Note that as calculated, the electron and positron when created have no extra energy, i.e. no velocity. To give them velocity we'd have to set Planck's photon energy equation equal to the full expression for Einstein's mass-energy equivalency:\[E ^{2}=\sqrt{m ^{2}c ^{4}+p ^{2}c ^{2}}\]where the p is the electron's (or positron's) momentum.

OpenStudy (anonymous):

I'm a full-pn physics geek, so the idea that someone was able to actually design a real experiment, one that could actually produce matter from light, pushed all of my physics buttons. Hopefully other physicists were as excited about this as me, and hopefully the announcement blew the minds of curious folk around the world.

OpenStudy (anonymous):

I'll turn my physics geek beacon off now.